Suppose an economy has 10,000 unemployed and searching for jobs and 90,000 people who are working. What is its unemployment rate? Now suppose 4,000 people looking for work get discouraged and give up their searches. What happenes to the unemployment rate? Would you interpret this as a good news for the economy ora bad news? Explain.
1
Expert's answer
2021-07-21T08:50:32-0400

The unemployment rate is:

U = 10,000/(90,000 + 10,000)×100% = 10%.

4,000 people looking for work get discouraged and give up their searches, then:

U = 6,000/(90,000 + 6,000)×100% = 6.25%.

So, the unemployment rate decreases, but it is a bad news for the economy, because the size of labour force also decreases.